An electric motor develops 65 kw of power as it lifts a loaded elevator 17.5 m in 35.0 s. How much force does the motor exert ?
1
Expert's answer
2019-01-10T15:42:22-0500

We can find how much force does the motor exert from the definition of the power:

"P = \frac{W}{t},""P = \frac{Fd}{t}."

Then, from this formula we can find the force that the motor exert:

"F = \frac{Pt}{d}."

Let's substitute the numbers:

"F = \frac{65 \cdot 10^3 W \cdot 35 s}{17.5 m} = 1.3 \cdot 10^5 N."

Answer:

"F= 1.3 \cdot 10^5 N"

.